OpenStudy (anonymous):

Find the linear and power function of lines passing through the points: (-3,19683) and (4,-147456)

OpenStudy (zzr0ck3r):

\(y=mx+b\) gives a "linear" curve. Use the points to figure out what \(m\) and \(b\) are. Do you know how to do this?

OpenStudy (zzr0ck3r):

What is the slope \(m\) of the line passing through the two points?

OpenStudy (anonymous):

yes y2-y1/x2-x1

OpenStudy (anonymous):

Im having trouble with the power function part

OpenStudy (zzr0ck3r):

what is the form of a power function?

OpenStudy (anonymous):

I don't know...I wasn't in lecture when it was taught..

OpenStudy (zzr0ck3r):

The form is \(y=ax^b\), so we need to figure out what \(a\) and \(b\) are. For this we do something similar to finding the solution to a system of linear equations (the addition method) but its more like a division method use the two points and we get \(19683=a(-3)^b\) and \(-147456=a(4)^b\) Do you see how I came up with that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

how do we find b?

OpenStudy (zzr0ck3r):

solve for a in one equation, then plug that back in for a in the other equation

OpenStudy (zzr0ck3r):

make sense?